Haswell Enterprises' bonds have a 10-year maturity, a 6.25% semiannual coupon, and a par value of $1,000. The going interest rate (rd) is 4%, based on semiannual compounding. What is the bond's price?

Select the correct answer.

a. $1,181.58
b. $1,191.06
c. $1,188.69
d. $1,186.32
e. $1,183.95
Answer #1

Coupon = (0.0625 * 1,000) / 2 = 31.25

Number of periods = 10 * 2 = 20

Rate = 4% / 2 = 2%

Price of Bond = Coupon * [ 1 - 1 / ( 1 + r)n] / r + FV / ( 1 + r)n

Price of Bond = 31.25 * [ 1 - 1 / ( 1 + 0.02)20] / 0.02 + 1000 / ( 1 + 0.02)20

Price of Bond = 31.25 * 16.351433 + 672.971333

Price of Bond = $1,183.95

Keys to use in a financial calculator:

2nd I/Y 2

FV 1000

PMT 31.25

N 20

I/Y 4

CPT PV